OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

OpenStudy (anonymous):

you need to put your equation in y = mx + b form. Can you do that ?

OpenStudy (anonymous):

@Devind23

OpenStudy (anonymous):

y=3x+3

OpenStudy (anonymous):

parallel to this line means same slope

OpenStudy (anonymous):

slope=3

OpenStudy (anonymous):

so by point slope form,the eqn. is (y--2)=3(x-4) ----->y+2=3x-12

OpenStudy (anonymous):

y-3x=-14

OpenStudy (anonymous):

or y=3x-14

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are welcome:D

OpenStudy (anonymous):

Can you help me out with my next 2 questions? They're the same just different numbers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

here slope=2....the slope of the perpendicular line should be negative reciprocal of this..so our line has a slope -1/2...(y-3)=(-1/2)(x-6)

OpenStudy (anonymous):

jus expand it

OpenStudy (anonymous):

y+1/2 x=6

OpenStudy (anonymous):

So it's Y= -1/2 + 6?

OpenStudy (anonymous):

-1/2 x

OpenStudy (anonymous):

you forgot the x

OpenStudy (anonymous):

Okay got it. Next question

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = x − 10 y = x + 10 y = −x − 10 y = −x + 10

OpenStudy (anonymous):

divide this line eqn by 3 so we get y= (2/3 x)-1

OpenStudy (anonymous):

slope of our line will be -3/2 ......(y-2)= -3/2 (x+8)

OpenStudy (anonymous):

I am getting y=(-3/2)x -10..but

OpenStudy (anonymous):

So the first one?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

its not in the options

OpenStudy (anonymous):

Oh the fractions from the questions didn't come up its there

OpenStudy (anonymous):

is it -3/2

OpenStudy (anonymous):

Okay last one

OpenStudy (anonymous):

2x - 3y = 3 -3y = -2x + 3 y = -2x/-3 + 3/-3 y = 2/3x - 1 perpendicular is negative reciprocal slope will be -3/2 y - y1 = m(x - x1).........(-8,2) y - 2 = -3/2(x -(-8) y - 2 = -3/2x + 8 y = -3/2x + 10

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = x + 7 y = −3x + 3 y = x + 3 y = −3x + 7

OpenStudy (anonymous):

1/3x is in the first and third answer

OpenStudy (anonymous):

slope =-3...(y-6)=-3(x+1) --->y=-3x+3

OpenStudy (anonymous):

Thank you so much! Dude whats your number just in case I have more trouble lol

OpenStudy (anonymous):

I will be here..

OpenStudy (anonymous):

or mail to dasuttan03@gmail.com

OpenStudy (anonymous):

okay..?

OpenStudy (anonymous):

Yeah